OpenStudy (anonymous):

Convert the given exponential function to the form indicated. Round all coefficients to four significant places. f(t) = 2.8(1.005)t; f(t) = Q0ekt

OpenStudy (amistre64):

you might be missing some notation in that

OpenStudy (amistre64):

\[1.05^t=e^{kt}\] \[t~ln(1.05)=kt\] \[ln(1.05)=k\]

OpenStudy (anonymous):

f(t) = 2.8(1.005)^t; f(t) = Q0e^kt

OpenStudy (amistre64):

well, ln(1.005) = k that is

OpenStudy (anonymous):

what is Q0?

OpenStudy (amistre64):

\[Q_o = 2.8\]

OpenStudy (anonymous):

Thank you so much

OpenStudy (amistre64):

youre welcome